OpenStudy (anonymous):

in a geometric sequence, t4 = 5x - 11, t5 = x + 5, and t6 = x - 1. Find the sum(s) of the infinite series if it (or they) exist(s).

ganeshie8 (ganeshie8):

the sum exists if |r| < 1

ganeshie8 (ganeshie8):

find the value of `x` that gives you |r| < 1

OpenStudy (anonymous):

but how do i find x?

ganeshie8 (ganeshie8):

notice thatt the given terms are consecutive terms in a geometric sequence, so the ratio of any two adjacent terms equals the common ratio \(r\) : \[\large r = \dfrac{x+5}{5x-11} = \dfrac{x-1}{x+5}\]

ganeshie8 (ganeshie8):

take the last two ratios and solve x : \[\large (x+5)^2 = (5x-11)(x-1)\]

OpenStudy (kropot72):

Assume that r = 1/2 Then we can write: \[\frac{1}{2}(5x-11)=x+5\] 2.5x - 5.5 = x + 5 1.5x = 10.5 x = 7 When x = 7 the geometric sequence becomes: 24, 12, 6, ...............

OpenStudy (kropot72):

I cannot get the method of @ganeshie8 to work.

OpenStudy (anonymous):

ahhh omg you're a lifesaver xD THANK YOU SO MUCH :D

OpenStudy (kropot72):

You're welcome :)

OpenStudy (kropot72):

My bad. The method of @ganeshie8 works perfectly, and should get higher marks!

OpenStudy (anonymous):

so either way, they both work?

OpenStudy (kropot72):

Yes, they both work.